how long does it take for the hoop to roll down without sliding from the inclined plane of length 2 m and height 10 cm ?
1
Expert's answer
2012-12-26T07:47:57-0500
L= 2 m - length and h=0.1 m - height.
Angle betweendirection of acceleration of gravitation -g and direction of motion of hoop is
alpha.
Find projection of g - a=g*cos(alpha), cos(alpha) equal h/L.cos meanscosine.
L=a*t^2/2, t- time of motion.
Find t. t=sqrt((2*L)/a)=sqrt((2*L)/(g*h/L)=sqrt((2*L^2)/(g*h))=2.857seconds.g=9.8 m/s^2